How much faster is Nitrogen gas vs Oxygen gas
Use three significant figures
1
Expert's answer
2020-04-28T12:21:27-0400

By Graham's Law of Diffusion,

Rate of diffusion is inversely proportional to the square root of the molar mass of gas.

"r\propto\frac{1}{\sqrt{M}}"

"M_{O_2}=32 \ g,M_{N_2}=28\ g."

"\implies \frac{r_{N_2}}{r_{O_2}}=\sqrt{\frac{M_{O_2}}{M_{N_2}}}=\sqrt{\frac{32}{28}}=1.069=1.07" (up to "3" significant digits)

"\implies" "r_{N_2}=1.07\times r_{O_2}" .

So,Nitrogen gas is "1.07" times faster than oxygen gas.